OpenStudy (anonymous):

Help Please! Solve for x -ax +4b>9

OpenStudy (anonymous):

\[ x <\frac{ 9-4b }{a }\]

OpenStudy (gorv):

take 4b on otherside

OpenStudy (anonymous):

\[x <\frac{ -4b+9 }{ a }\]

OpenStudy (anonymous):

\[x > \frac{ 4b+9 }{ -a }\]

OpenStudy (anonymous):

these are the choices. So how do I solve for x here?

OpenStudy (anonymous):

can you show me the steps?

OpenStudy (anonymous):

like this @gorv ?

OpenStudy (anonymous):

so what would be the next step?

OpenStudy (anonymous):

-ax +4b-4b> 9 - 4b

OpenStudy (anonymous):

@lalaly can you help me to understand the process

OpenStudy (gorv):

-ax<9-4b

OpenStudy (gorv):

now ax ix negative ........so make it positive...by multiplying wiith - and due to this inequality also changes

OpenStudy (gorv):

ax>-(9-4b) now move a on other side that"s it

OpenStudy (anonymous):

I am still bit confuse here. Do we need to divide both side or multiplying ? I thought divide to make x alone

OpenStudy (anonymous):

did you mean like this ? -ax*ax >- (9-4b)

OpenStudy (anonymous):

@wolfe8

OpenStudy (anonymous):

can you help me bit with this ?

OpenStudy (anonymous):

@terenzreignz

OpenStudy (wolfe8):

Remember that you are dividing by -a

OpenStudy (anonymous):

-ax/-a > 9-4b/-a x > 9-4b/a like this?

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@ajprincess can you help me with this?

OpenStudy (mathmale):

Solve for x -ax +4b>9 We want to isolate x. One of the easier ways of doing this is to make the ax term positive. We can do that by adding ax to both sides, as follows: -ax + 4b > 9 +ax +ax ----------------- 4b > 9 + ax To isolate the ax term, subtract 9 from both sides of this inequality: 4b > 9 + ax -9 -9 ------- ------ 4b-9 > ax Finally, isolate x by dividing both sides of the inequality by a: 4b-9 > ax ------ --- a a Thus, \[\frac{ 4b-9 }{ a }>x\],

OpenStudy (mathmale):

or\[x<\frac{ 4b-9 }{ a }\]

OpenStudy (anonymous):

thanks so much @mathmale helping me to understand this!!!!

OpenStudy (mathmale):

Very happy to do it! See you again.

OpenStudy (anonymous):

:) thanks alot again!!!

OpenStudy (anonymous):

Thanks so much all of you for coming here and trying hard to help me @wolfe8 @gorv!!!!! And I am really sorry for troubling all of you

OpenStudy (anonymous):

Thank you @ajprincess coming here for help :)

OpenStudy (anonymous):

There is problem it is not truly correct I got this answer earlier and it was wrong @mathmale :(

OpenStudy (anonymous):

I did not noticed when you're showing step

OpenStudy (mathmale):

would you please go back to the very beginning and make sure, absolutely sure, that you've copied the homework problem correctly?

OpenStudy (anonymous):

yep i did correctly

OpenStudy (anonymous):

@satellite73

OpenStudy (mathmale):

Let's try a different method and see whether that produces a different result or not. There can be only one correct result, and that result should be independent of the method used to obtain it.

OpenStudy (mathmale):

Solve for x: -ax + 4b > 9 Isolate -ax by subtracting 4b from both sides: (you do it, please)

OpenStudy (anonymous):

-ax +4b-4b > 9 - 4b like this?

OpenStudy (mathmale):

Yes. that reduces to -ax > 9-4b. Our next task is to isolate x. To do this, we must divide both sides of the inequality above by -a AND we must reverse the direction of the inequality sign (because of dividing by a (-) number). -ax > 9-4b --- ----- -a -a 9-4b results in x < ---------- -a and this is equivalent to 4b-9 x<------- a

OpenStudy (anonymous):

thanks so much for help and I am really for bother you again and again. Please forgive me

OpenStudy (mathmale):

Again, I'm very happy to be of help, especially when it leads to your understanding of the problem we've been working on. :)

OpenStudy (anonymous):

thanks so much for help again and tolerating me, haha :)

OpenStudy (wolfe8):

Yeah my bad I fell asleep. I meant to say you got it right. Good luck.

OpenStudy (anonymous):

No you did amazing, you worked really hard. It was me who is stupid.